hugo paid 95.89 for a new pair of running shoes at the check-out register, he was charged a additional 8% what was the total costof the running shoes?
Answer by fcabanski(1391)

When determining interest on a sale the formula is P = S(1+i) where P is the final price, S is the marked sales price and i is the interest or tax.


The wording of this problem isn't clear. If he paid $95.89 including the 8% (.08) tax and the problem asks for the marked price (before tax) then it's:


95.89 = S(1.08)
S = 95.89/1.08 = $88.79


If he paid 95.89 (the marked price) plus 8% (.08) tax then it's:


P = 95.89(1.08)=$103.56

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com